# A Simple yet Effective Test-Time Adaptation for Zero-Shot Monocular Metric Depth Estimation

### v2

@@ Abstract
The recent development of foundation models for monocular depth estimation such as Depth Anything paved the way to zero-shot monocular depth estimation. Since it returns an affine-invariant disparity map, the favored technique to recover the metric depth consists in fine-tuning the model. However, this stage is not straightforward, it can be costly and time-consuming because of the training and the creation of the dataset. The latter must contain images captured by the camera that will be used at test time and the corresponding ground truth. Moreover, the fine-tuning may also degrade the generalizing capacity of the original model. Instead, we propose in this paper a new method to rescale Depth Anything predictions using 3D points provided by sensors or techniques such as low-resolution LiDAR or structure-from-motion with poses given by an IMU. This approach avoids fine-tuning and preserves the generalizing power of the original depth estimation model while being robust to the noise of the sparse depth or of the depth model. Our experiments highlight enhancements relative to zero-shot monocular metric depth estimation methods, competitive results compared to fine-tuned approaches and a better robustness than depth completion approaches. Code available at gitlab.ensta.fr/ssh/monocular-depth-rescaling.
@@ I Introduction
Despite the growth of 3D perception sensors such as LiDAR, time-of-flight, structured light or stereo cameras, for robotic systems, traditional monocular cameras remain a key sensor for any robot setup. In addition to being a cheaper solution, monocular depth estimation can also offer denser outputs as well as a larger depth range. The increase in the number of open datasets for monocular depth estimation and the development of neural network architectures such as Vision Transformers that scale well with the size of the training dataset [1] allowed the emergence of foundation models for monocular depth estimation [2].
@@
Predicting metric depth (also referred to as absolute depth) from a single image is fundamentally an impossible task due to scaling ambiguities. However, monocular depth estimation methods [3] can achieve very good performances in a defined context. These approaches have learned depth cues in images sampled from a certain distribution of environments that have been captured by a camera with fixed calibration. Thus, such models, trained on a single dataset, generalize poorly to images taken with a different camera. To avoid this issue during training on multiple datasets, some methods [4, 2, 5] learn an affine-invariant depth or disparity which allows impressive results on zero-shot relative depth estimation benchmarks. Then, they propose to fine-tune their pre-trained models on the target domain to make metric depth predictions, i.e., on a dataset composed of images captured with the target camera calibration. Such a solution is costly in real cases due to the creation of the dataset and the training computation. Moreover, it must be performed for each new camera calibration. Several solutions have been proposed to solve this issue by explicitly taking into account the camera calibration in the method [6] or trying to learn it [7]. Depth completion methods propose another alternative [8, 9] as they take as input some sparse depth measurements. However, all these approaches may be more costly at inference or cannot be trained on image datasets with unknown calibration or unknown ground truth depth such as ImageNet [10] contrary to methods like Depth Anything [2, 5].
@@
In this paper, following depth completion approaches, we investigate a test-time adaptation that leverages sparse depth measurements for solving the scale ambiguity in order to perform zero-shot monocular metric depth estimation given affine-invariant disparity predictions. Thus, an additional sensor is used to obtain some reference 3D points that are exploited to recover the scaling parameters. For the sake of brevity, we will refer to rescaling the process of finding an affine transformation and applying it to recover the metric depth. We focus our study on sparse depth that can be provided by other sources including low-resolution LiDARs (with 16 and 32 beams), 2D LiDARs (with a single beam) that are often used for indoor robotics and structure-from-motion. In the latter, we assume a metric relative camera pose is given by an IMU. The advantage of our approach is twofold. On the one hand, it can be used with any monocular depth estimation model such as Depth Anything V1 and V2 [2, 5] with a high generalization ability due to their large and diverse training dataset. On the other hand, our method does not require any costly fine-tuning on the target domain and provides instant adaptation. We conducted extensive experiments to evaluate our approaches on standard metric depth estimation benchmarks and demonstrate robustness to a noisy sparse depth or to a drop in sparse depth density.
@@ II Related work
Figure 1: Illustration of our method. First, an affine-invariant disparity map $d$ is predicted from an image with a neural network such as Depth Anything [2] while in parallel a sensor is used to estimate a set of 3D points $P$. $P$ and the corresponding values in $d$ are then used to estimate the scaling parameters $\alpha$ and $\beta$ using a RANSAC [11]. The parameters are applied to $d$ to recover the metric depth $D$.
@@ II-A Monocular depth estimation
While pioneer works on monocular depth estimation relied on Markov Random Fields [12, 13], subsequent ones have shown the effectiveness of convolutional neural networks [14, 15] then transformers for this task [16]. More recently, benefiting from advances in image generation [17], impressive results have been obtained using diffusion models [18, 19]. Regarding the output, monocular depth estimation has been initially addressed as a regression problem [14], before moving to classification approaches with discrete bins that show better performances [3, 20]. Monocular depth estimation methods can also be divided into whether they are supervised [14, 15, 3] or self-supervised [21]. Recently, the multiplication of depth estimation benchmark allowed training models on multiple domains at once [4] and paved the way to zero-shot monocular depth estimation. Due to the inherent scale ambiguity of depth estimation from a single image, these methods are mostly trained to produce affine-invariant depth or disparity predictions [4, 16, 2]. Our work aims to be used with any depth estimation model as long as it returns disparity maps that are accurate within an affine transformation whatever its other characteristics in terms of training set or architecture.
@@ II-B Scale estimation for monocular depth
A major issue of monocular depth prediction relies on scale ambiguity which means that the true size of an object cannot be recovered for sure from a single image. Since most monocular depth estimation models are trained for a specific camera calibration, using them with another camera leads to ill-scaled predictions. The most common way to recover the metric depth in this situation is to fine-tune the model on a dataset collected with the camera that will be used at inference [4, 16]. In practice, this solution is costly to implement as it requires the creation of an image dataset with the relative ground truth and a new training. Other works focus on the temporal consistency of the scale of depth predictions [22, 23]. Closer to our work, [24] learns to predict scaled depth maps from affine-invariant disparity maps and visual-inertial odometry. Also, there exists an extensive literature on depth completion which studies neural network architectures that take as inputs both an image and a sparse depth map [25, 8, 9]. In [26], authors propose to estimate the scale factor for a target domain from a model that has been jointly trained on a source domain with known ground truth and on a target domain without depth annotation. In contrast, we propose to rescale at test time any disparity prediction that is correct up to an affine transformation with no additional training or fine-tuning but by leveraging reference 3D points provided by an external sensor or technique. Furthermore, relying on external sensor makes our approach adaptable to any camera calibration.
@@ II-C Zero-shot monocular metric depth estimation
ZoeDepth [27] is the first zero-shot monocular metric depth estimation method. It first consists in relative depth pre-training of the MiDaS [4] backbone then fine-tuning two metric bin modules, one for indoor scenes and the other for outdoor scenes. More recently, other methods have been proposed without any fine-tuning. Thus, ScaleDepth [28] decomposes metric depth estimation in relative depth estimation and scale estimation each with a dedicated module. It can also leverage a text description of the scene to guide the supervision. In [6], authors introduce a dedicated architecture that takes as input the calibration matrix in addition to the image. On the contrary, [29, 30] make predictions from the images only but applies transformation on the input images so the predictions are invariant to the image size or the camera calibration. UniDepth [7] estimates an internal representation of the camera calibration directly from the input images only. These approaches still have drawbacks since they are often more costly at inference or need the calibration of the images even at the training stage. This prevents exploiting image datasets with unknown calibration unlike Depth Anything V1 [2] and V2 [5] which leveraged a distillation strategy with such datasets to improve their generalization abilities.
@@ III Method
Let $\phi$ be a monocular depth estimation model such as [4, 2] that is trained to predict an affine-invariant disparity map $d\in\mathbb{R}^{H\times W}$ given an input RGB image $I\in\mathbb{R}^{H\times W\times 3}$ where $H$ and $W$ are the height and the width of the image $I$. Therefore, the metric or absolute depth map $D\in\mathbb{R}^{H\times W}$ that corresponds to the inverse of the metric or absolute disparity $D^{-1}$ is given by the relation:
@@
Our method, illustrated in Fig. 1, aims at recovering the metric depth map $D$ at test time from the affine-invariant disparity map $d$ and a set of $N$ reference 3D points $P\in\mathbb{R}^{N\times 3}$ by regressing the parameters $\alpha$ and $\beta$. First, we perform a bilinear sampling on the affine-invariant map $d$ at the locations of the projection of the reference 3D points $P$ on the image plan so as to have an affine-invariant disparity value for each reference 3D point. Second, we leverage linear regression to estimate $\alpha$ and $\beta$ parameters. To favor the robustness against potential outliers in the reference 3D points, we use a RANSAC algorithm [11], but other robust regression methods could be considered.
@@
As our method relies on affine-invariant disparity maps, we can exploit, without any fine-tuning, foundation models like Depth Anything V1 or V2 [2, 5] that are trained to predict such outputs. Furthermore, the robustness to outliers enabled by the RANSAC allows low-quality sensors to be used to obtain the reference 3D points. Thus, we can provide metric depth maps whatever the environment or the camera with no fine-tuning. Nevertheless, since Depth Anything processes each image independently and is supervised in such a way there is no guarantee its disparity predictions are normalized identically even for consecutive images of a video, the scaling factor $\alpha$ and the offset $\beta$ must be estimated for each image.
@@ IV-A Dataset and metrics
To evaluate performance in zero-shot monocular metric depth estimation, we focus on standard monocular depth estimation benchmarks that have not been used to pretrain Depth Anything [2]. This includes indoor datasets: NYUv2 [31], SUN-RGBD [32], IBIMS-1 [33] and DIODE indoor [34] and outdoor datasets: KITTI [35] , DDAD [36] and DIODE outdoor [34]. We adopt standard depth estimation metrics (see [14, 2]) to compare our method to other approaches of the literature:
@@
where $\Omega$ is the set of pixels for which the ground truth is available and |.||.|| . | applied to a set returns its cardinal and the absolute value of a scalar otherwise. For a pixel $p\in\Omega$, $\hat{D}(p)$ is the estimation corresponding the ground truth depth $D^{*}(p)$.
@@ IV-B Implementation details
All experiments have been conducted using Depth Anything V1 [2] with ViT Large [1] without fine-tuning unless otherwise mentioned. We adopt Depth Anything V1 [2] code base and settings for evaluation except for comparison with depth completion methods as detailed later. We simulate low-resolution LiDARs by evenly selecting as many horizontal lines in the ground truth depth maps as the number of beams in 32-laser, 16-laser or 2D LiDARs (i.e., with a single beam). In contrast, no ground truth depth is used as input when rescaling with SFM. We study rescaling with SFM for KITTI [35] and DDAD datasets [36] only as they both consist of temporal sequences of images including the pose estimations. To obtain the reference 3D points, we first extract matching keypoints in the target image and the previous one using SIFT [37] or OmniGlue [38] and triangulate them using the pose between these images. Since this strategy requires enough displacement magnitude between the two images, we only consider image couples with a rotation higher than $5$ degrees or a translation greater than $1.5$ and $2$ meters for KITTI and DDAD datasets, respectively. We note that this threshold is the only hyperparameter that needs to be tuned in our approach.
@@ IV-C Comparison with monocular depth esimation methods
TABLE I: Quantitative results on the NYUv2 dataset [31] (indoor). (ZS) means Zeros-Shot, (FT) stands for Fine-Tuned on NYUv2.
Methods $\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$
ZeroDepth [6] (ZS) 0.926 0.081 0.338
Metric3D [29] (ZS) 0.944 0.083 0.310
Metric3D V2 [30] (ZS) 0.975 0.063 0.251
@@
ZeroDepth [6] (FT) 0.954 0.074 0.269
ZoeDepth [27] (FT) 0.955 0.075 0.270
Metric3D V2 [30] (FT) 0.989 0.047 0.183
ScaleDepth [28] (FT) 0,957 0,074 0,267
@@
TABLE II: Quantitative results on the KITTI dataset [35] (outdoor). (ZS) means Zeros-Shot, (FT) stands for Fine-Tuned on KITTI.
Methods $\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$
ZeroDepth [6] (ZS) 0.910 0.102 4.044
Metric3D [29] (ZS) 0.964 0.058 2.770
Metric3D V2 [30] (ZS) 0.974 0.052 2.511
ZoeDepth [27] (FT) 0.971 0.057 2.281
ZeroDepth [6] (FT) 0.968 0.053 2.087
Metric3D V2 [30] (FT) 0.985 0.044 1.985
ScaleDepth [28] (FT) 0,980 0,048 1,987
@@
Ours w/ SFM (SIFT [37]) 0.893 0.103 3.920
Ours w/ SFM (OmniGlue [38]) 0.925 0.093 3.562
TABLE III: Quantitative results on different zero-shot indoor benchmarks. FT NYUv2 stands for Fine-Tuned on NYUv2 dataset [31].
Methods SUN-RGBD [32] IBIMS-1 [33] DIODE Indoor [34]
$\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$
Metric3D [29] 0.144 0.252
Metric3D V2 [30] 0.185 0.592 0.093 0.389
Unidepth [7] 0.966 0.797 0.774
@@
Depth Anything [2] (FT NYUv2) 0.658 0.500 0.616 0.714 0.150 0.593 0.303 0.325 1.476
@@
TABLE IV: Quantitative results on different zero-shot outdoor benchmarks. FT KITTI stands for Fine-Tuned on KITTI dataset [35].
Method DIODE Outdoor [34] DDAD [36]
$\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$
ZoeDepth (FT KITTI) [27] 0.835 0.129 7.108
ZeroDepth [6] 0.814 0.156 10.678
@@
Metric3D V2 [30] 0.221 3.897
Unidepth [7] 0.864
ScaleDepth [28] (FT KITTI) 0.333 0.605 6.950 0.863 0.120 6.378
Depth Anything V1 [2] (FT KITTI) 0.288 0.794 6.641 0.886 0.105 5.931
Our w/ LiDAR 1 beam 0.689 0.880 6.222 0.706 0.326 9.229
@@
Our w/ SFM (SIFT) 0.776 0.161 5.929
Our w/ SFM (Omniglue) 0.947 0.112 5.300
In Tab. I, we compare our LiDAR-based rescaling approach with zero-shot monocular depth estimation methods and other depth estimation methods that have been fine-tuned on the NYUv2 dataset [31]. We conducted a similar study in Tab. II on the KITTI dataset [35] with in addition results of rescaling with 3D reference points provided by structure-from-motion. Tab. III and Tab. IV provide zero-shot performance on indoor and outdoor datasets, respectively. For each dataset, our results show rescaling using LiDARs with different numbers of beams. Additionally, we present results of rescaling with structure-from-motion for the DDAD dataset [36]. We do not provide results of rescaling with structure-from-motion on NYUv2 [31], SUN-RGBD [32], IBIMS-1 [33] and DIODE Indoor and Outdoor [34] because they do not consist of temporal image sequences with poses. The results of the monocular depth estimation methods we compare against were directly taken from their respective papers, which explain missing numbers when they have not been reported by their authors.
@@
Our experiments highlight an overall benefit of using our rescaling approach with $32$-laser or $16$-laser LiDAR rather than using zero-shot monocular metric depth estimation. Thus, we observe on average $46$% improvement on $\delta_{1}$, $0.5$% on AbsRel and $47$% on RMS relative to the other zero-shot methods. In contrast, the methods that have been fine-tuned on the same domain as the test set (see lines with (FT) in Tab. I and Tab. II) compare favorably to ours. These approaches benefit from an additional in-domain training which is often costly due to dataset creation and training computation. We notice that our method is sensitive to the domain with on average $28$% and $13$% enhancements on indoor (including NYUv2) and outdoor (including KITTI) datasets, respectively, relative to other zero-shot methods whatever the metric. Regarding performance with 2D LiDAR, the same sensitivity to the domain is apparent. However, the results are a bit lower than zero-shot and fine-tuned metric depth estimation methods and lower than our other approaches, especially on outdoor datasets while remaining competitive on indoor datasets. Interestingly, we observe that increasing the number of beams from 16 to 32 does not necessarily improve the performance.

### v3

@@ Abstract
The recent development of foundation models for monocular depth estimation such as Depth Anything paved the way to zero-shot monocular depth estimation. Since it returns an affine-invariant disparity map, the favored technique to recover the metric depth consists in fine-tuning the model. However, this stage is not straightforward, it can be costly and time-consuming because of the training and the creation of the dataset. The latter must contain images captured by the camera that will be used at test time and the corresponding ground truth. Moreover, the fine-tuning may also degrade the generalizing capacity of the original model. Instead, we propose in this paper a new method to rescale Depth Anything predictions using 3D points provided by sensors or techniques such as low-resolution LiDAR or structure-from-motion with poses given by an IMU. This approach avoids fine-tuning and preserves the generalizing power of the original depth estimation model while being robust to the noise of the sparse depth, of the camera-LiDAR calibration or of the depth model. Our experiments highlight enhancements relative to zero-shot monocular metric depth estimation methods, competitive results compared to fine-tuned approaches and a better robustness than depth completion approaches. Code available at github.com/ENSTA-U2IS-AI/depth-rescaling.
@@ I Introduction
Despite the growth of 3D perception sensors such as LiDAR, time-of-flight, structured light or stereo cameras, for robotic systems, traditional monocular cameras remain a key sensor for any robot setup. In addition to being a cheaper solution, monocular depth estimation can also offer denser outputs as well as a larger depth range. The increase in the number of open datasets for monocular depth estimation and the development of neural network architectures such as Vision Transformers that scale well with the size of the training dataset [dosovitskiy2020image] allowed the emergence of foundation models for monocular depth estimation [yang2024depth].
@@
Predicting metric depth (also referred to as absolute depth) from a single image is fundamentally an impossible task due to scaling ambiguities. However, monocular depth estimation methods [bhat2021adabins] can achieve very good performances in a defined context. These approaches have learned depth cues in images sampled from a certain distribution of environments that have been captured by a camera with fixed calibration. Thus, such models, trained on a single dataset, generalize poorly to images taken with a different camera. To avoid this issue during training on multiple datasets, some methods [ranftl2020towards, yang2024depth, yang2024depthv2] learn an affine-invariant depth or disparity which allows impressive results on zero-shot relative depth estimation benchmarks. Then, they propose to fine-tune their pre-trained models on the target domain to make metric depth predictions, i.e., on a dataset composed of images captured with the target camera calibration. Such a solution is costly in real cases due to the creation of the dataset and the training computation. Moreover, it must be performed for each new camera calibration. Several solutions have been proposed to solve this issue by explicitly taking into account the camera calibration in the method [guizilini2023towards] or trying to learn it [piccinelli2024unidepth]. Depth completion methods propose another alternative [zhang2023completionformer, tang2024bilateral] as they take as input some sparse depth measurements. However, all these approaches may be more costly at inference or cannot be trained on image datasets with unknown calibration or unknown ground truth depth such as ImageNet [deng2009imagenet] contrary to methods like Depth Anything [yang2024depth, yang2024depthv2].
@@
In this paper, following depth completion approaches, we investigate a test-time adaptation that leverages sparse depth measurements for solving the scale ambiguity in order to perform zero-shot monocular metric depth estimation given affine-invariant disparity predictions. Thus, an additional sensor is used to obtain some reference 3D points that are exploited to recover the scaling parameters. For the sake of brevity, we will refer to rescaling the process of finding an affine transformation and applying it to recover the metric depth. We focus our study on sparse depth that can be provided by other sources including low-resolution LiDARs (with 16 and 32 beams), 2D LiDARs (with a single beam) that are often used for indoor robotics and structure-from-motion. In the latter, we assume a metric relative camera pose is given by an IMU. The advantage of our approach is twofold. On the one hand, it can be used with any monocular depth estimation model such as Depth Anything V1 and V2 [yang2024depth, yang2024depthv2] with a high generalization ability due to their large and diverse training dataset. On the other hand, our method does not require any costly fine-tuning on the target domain and provides instant adaptation. We conducted extensive experiments to evaluate our approaches on standard metric depth estimation benchmarks and demonstrate robustness to a noisy sparse depth, to errors in camera-LiDAR calibration or to a drop in sparse depth density.
@@ II Related work
Figure 1: Illustration of our method. First, an affine-invariant disparity map $d$ is predicted from an image with a neural network such as Depth Anything [yang2024depth] while in parallel a sensor is used to estimate a set of 3D points $P$. $P$ and the corresponding values in $d$ are then used to estimate the scaling parameters $\alpha$ and $\beta$ using a RANSAC [fischler1981random]. The parameters are applied to $d$ to recover the metric depth $D$.
@@ II-A Monocular depth estimation
While pioneer works on monocular depth estimation relied on Markov Random Fields [saxena2005learning], subsequent ones have shown the effectiveness of convolutional neural networks [eigen2014depth, laina2016deeper] then transformers for this task [ranftl2021vision]. More recently, benefiting from advances in image generation [rombach2022high], impressive results have been obtained using diffusion models [saxena2024surprising, ke2024repurposing]. Regarding the output, monocular depth estimation has been initially addressed as a regression problem [eigen2014depth], before moving to classification approaches with discrete bins that show better performances [bhat2021adabins, bhat2022localbins]. Monocular depth estimation methods can also be divided into whether they are supervised [eigen2014depth, laina2016deeper, bhat2021adabins] or self-supervised [godard2019digging]. Recently, the multiplication of depth estimation benchmark allowed training models on multiple domains at once [ranftl2020towards] and paved the way to zero-shot monocular depth estimation. Due to the inherent scale ambiguity of depth estimation from a single image, these methods are mostly trained to produce affine-invariant depth or disparity predictions [ranftl2020towards, ranftl2021vision, yang2024depth]. Our work aims to be used with any depth estimation model as long as it returns disparity maps that are accurate within an affine transformation whatever its other characteristics in terms of training set or architecture.
@@ II-B Scale estimation for monocular depth
A major issue of monocular depth prediction relies on scale ambiguity which means that the true size of an object cannot be recovered for sure from a single image. Since most monocular depth estimation models are trained for a specific camera calibration, using them with another camera leads to ill-scaled predictions. The most common way to recover the metric depth in this situation is to fine-tune the model on a dataset collected with the camera that will be used at inference [ranftl2020towards, ranftl2021vision]. In practice, this solution is costly to implement as it requires the creation of an image dataset with the relative ground truth and a new training. Other works focus on the temporal consistency of the scale of depth predictions [zhang2019exploiting, li2021enforcing]. Closer to our work, [wofk2023monocular] learns to predict scaled depth maps from affine-invariant disparity maps and visual-inertial odometry. Also, there exists an extensive literature on depth completion which studies neural network architectures that take as inputs both an image and a sparse depth map [ma2018sparse, zhang2023completionformer, tang2024bilateral]. In [dana2024more], authors propose to estimate the scale factor for a target domain from a model that has been jointly trained on a source domain with known ground truth and on a target domain without depth annotation. In contrast, we propose to rescale at test time any disparity prediction that is correct up to an affine transformation with no additional training or fine-tuning but by leveraging reference 3D points provided by an external sensor or technique. Furthermore, relying on external sensor makes our approach adaptable to any camera calibration.
@@ II-C Zero-shot monocular metric depth estimation
ZoeDepth [bhat2023zoedepth] is the first zero-shot monocular metric depth estimation method. It first consists in relative depth pre-training of the MiDaS [ranftl2020towards] backbone then fine-tuning two metric bin modules, one for indoor scenes and the other for outdoor scenes. More recently, other methods have been proposed without any fine-tuning. Thus, ScaleDepth [zhu2024scaledepth] decomposes metric depth estimation in relative depth estimation and scale estimation each with a dedicated module. It can also leverage a text description of the scene to guide the supervision. In [guizilini2023towards], authors introduce a dedicated architecture that takes as input the calibration matrix in addition to the image. On the contrary, [yin2023metric3d, hu2024metric3d] make predictions from the images only but applies transformation on the input images so the predictions are invariant to the image size or the camera calibration. UniDepth [piccinelli2024unidepth] estimates an internal representation of the camera calibration directly from the input images only. These approaches still have drawbacks since they are often more costly at inference or need the calibration of the images even at the training stage. This prevents exploiting image datasets with unknown calibration unlike Depth Anything V1 [yang2024depth] and V2 [yang2024depthv2] which leveraged a distillation strategy with such datasets to improve their generalization abilities.
@@ III Method
Let $\phi$ be a monocular depth estimation model such as [ranftl2020towards, yang2024depth] that is trained to predict an affine-invariant disparity map $d\in\mathbb{R}^{H\times W}$ given an input RGB image $I\in\mathbb{R}^{H\times W\times 3}$ where $H$ and $W$ are the height and the width of the image $I$. Therefore, the metric or absolute depth map $D\in\mathbb{R}^{H\times W}$ that corresponds to the inverse of the metric or absolute disparity $D^{-1}$ is given by the relation:
@@
Our method, illustrated in Fig. 1, aims at recovering the metric depth map $D$ at test time from the affine-invariant disparity map $d$ and a set of $N$ reference 3D points $P\in\mathbb{R}^{N\times 3}$ by regressing the parameters $\alpha$ and $\beta$. First, we perform a bilinear sampling on the affine-invariant map $d$ at the locations of the projection of the reference 3D points $P$ on the image plan so as to have an affine-invariant disparity value for each reference 3D point. Second, we leverage linear regression to estimate $\alpha$ and $\beta$ parameters. To favor the robustness against potential outliers in the reference 3D points, we use a RANSAC algorithm [fischler1981random], but other robust regression methods could be considered.
@@
As our method relies on affine-invariant disparity maps, we can exploit, without any fine-tuning, foundation models like Depth Anything V1 or V2 [yang2024depth, yang2024depthv2] that are trained to predict such outputs. Furthermore, the robustness to outliers enabled by the RANSAC allows low-quality sensors to be used to obtain the reference 3D points. Thus, we can provide metric depth maps whatever the environment or the camera with no fine-tuning. Nevertheless, since Depth Anything processes each image independently and is supervised in such a way there is no guarantee its disparity predictions are normalized identically even for consecutive images of a video, the scaling factor $\alpha$ and the offset $\beta$ must be estimated for each image.
@@ IV-A Dataset and metrics
To evaluate performance in zero-shot monocular metric depth estimation, we focus on standard monocular depth estimation benchmarks that have not been used to pretrain Depth Anything [yang2024depth]. This includes indoor datasets: NYUv2 [nyu], SUN-RGBD [song2015sun], IBIMS-1 [ibims] and DIODE indoor [diode_dataset] and outdoor datasets: KITTI [kitti] , DDAD [guizilini20203d] and DIODE outdoor [diode_dataset]. We adopt standard depth estimation metrics (see [eigen2014depth, yang2024depth]) to compare our method to other approaches of the literature:
@@
where $\Omega$ is the set of pixels for which the ground truth is available and $|.|$ applied to a set returns its cardinal and the absolute value of a scalar otherwise. For a pixel $p\in\Omega$, $\hat{D}(p)$ is the estimation corresponding the ground truth depth $D^{*}(p)$.
@@ IV-B Implementation details
All experiments have been conducted using Depth Anything V1 [yang2024depth] with ViT Large [dosovitskiy2020image] without fine-tuning unless otherwise mentioned. We adopt Depth Anything V1 [yang2024depth] code base and settings for evaluation except for comparison with depth completion methods as detailed later. We simulate low-resolution LiDARs by evenly selecting as many horizontal lines in the ground truth depth maps as the number of beams in 32-laser, 16-laser or 2D LiDARs (i.e., with a single beam). In contrast, no ground truth depth is used as input when rescaling with SFM. We study rescaling with SFM for KITTI [kitti] and DDAD datasets [guizilini20203d] only as they both consist of temporal sequences of images including the pose estimations. To obtain the reference 3D points, we first extract matching keypoints in the target image and the previous one using SIFT [lowe1999object] or OmniGlue [jiang2024omniglue] and triangulate them using the pose between these images. Since this strategy requires enough displacement magnitude between the two images, we only consider image couples with a rotation higher than $5$ degrees or a translation greater than $1.5$ and $2$ meters for KITTI and DDAD datasets, respectively. We note that this threshold is the only hyperparameter that needs to be tuned in our approach.
@@ IV-C Comparison with monocular depth esimation methods
TABLE I: Quantitative results on the NYUv2 dataset [nyu] (indoor). (ZS) means Zeros-Shot, (FT) stands for Fine-Tuned on NYUv2.
Methods $\delta_{1}$ $\uparrow$ AbsRel $\downarrow$ RMS $\downarrow$
ZeroDepth [guizilini2023towards] (ZS) 0.926 0.081 0.338
Metric3D [yin2023metric3d] (ZS) 0.944 0.083 0.310
Metric3D V2 [hu2024metric3d] (ZS) 0.975 0.063 0.251
@@
ZeroDepth [guizilini2023towards] (FT) 0.954 0.074 0.269
ZoeDepth [bhat2023zoedepth] (FT) 0.955 0.075 0.270
Metric3D V2 [hu2024metric3d] (FT) 0.989 0.047 0.183
ScaleDepth [zhu2024scaledepth] (FT) 0,957 0,074 0,267
@@
TABLE II: Quantitative results on the KITTI dataset [kitti] (outdoor). (ZS) means Zeros-Shot, (FT) stands for Fine-Tuned on KITTI.
Methods $\delta_{1}$ $\uparrow$ AbsRel $\downarrow$ RMS $\downarrow$
ZeroDepth [guizilini2023towards] (ZS) 0.910 0.102 4.044
Metric3D [yin2023metric3d] (ZS) 0.964 0.058 2.770
Metric3D V2 [hu2024metric3d] (ZS) 0.974 0.052 2.511
ZoeDepth [bhat2023zoedepth] (FT) 0.971 0.057 2.281
ZeroDepth [guizilini2023towards] (FT) 0.968 0.053 2.087
Metric3D V2 [hu2024metric3d] (FT) 0.985 0.044 1.985
ScaleDepth [zhu2024scaledepth] (FT) 0,980 0,048 1,987
@@
Ours w/ SFM (SIFT [lowe1999object]) 0.893 0.103 3.920
Ours w/ SFM (OmniGlue [jiang2024omniglue]) 0.925 0.093 3.562
TABLE III: Quantitative results on different zero-shot indoor benchmarks. FT NYUv2 stands for Fine-Tuned on NYUv2 [nyu].
Methods SUN-RGBD [song2015sun] IBIMS-1 [ibims] DIODE Indoor [diode_dataset]
$\delta_{1}$ $\uparrow$ AbsRel $\downarrow$ RMS $\downarrow$ $\delta_{1}$ $\uparrow$ AbsRel $\downarrow$ RMS $\downarrow$ $\delta_{1}$ $\uparrow$ AbsRel $\downarrow$ RMS $\downarrow$
Metric3D [yin2023metric3d] 0.144 0.252
Metric3D V2 [hu2024metric3d] 0.185 0.592 0.093 0.389
Unidepth [piccinelli2024unidepth] 0.966 0.797 0.774
@@
Depth Anything [yang2024depth] (FT NYUv2) 0.658 0.500 0.616 0.714 0.150 0.593 0.303 0.325 1.476
@@
TABLE IV: Quantitative results on different zero-shot outdoor benchmarks. FT KITTI stands for Fine-Tuned on KITTI [kitti].
Methods DIODE Outdoor [diode_dataset] DDAD [guizilini20203d]
$\delta_{1}$ $\uparrow$ AbsRel $\downarrow$ RMS $\downarrow$ $\delta_{1}$ $\uparrow$ AbsRel $\downarrow$ RMS $\downarrow$
ZoeDepth (FT KITTI) [bhat2023zoedepth] 0.835 0.129 7.108
ZeroDepth [guizilini2023towards] 0.814 0.156 10.678
@@
Metric3D V2 [hu2024metric3d] 0.221 3.897
Unidepth [piccinelli2024unidepth] 0.864
ScaleDepth [zhu2024scaledepth] (FT KITTI) 0.333 0.605 6.950 0.863 0.120 6.378
Depth Anything V1 [yang2024depth] (FT KITTI) 0.288 0.794 6.641 0.886 0.105 5.931
Ours w/ LiDAR 1 beam 0.689 0.880 6.222 0.706 0.326 9.229
@@
Ours w/ SFM (SIFT) 0.776 0.161 5.929
Ours w/ SFM (Omniglue) 0.947 0.112 5.300
In Tab. I, we compare our LiDAR-based rescaling approach with zero-shot monocular depth estimation methods and other depth estimation methods that have been fine-tuned on the NYUv2 dataset [nyu]. We conducted a similar study in Tab. II on the KITTI dataset [kitti] with in addition results of rescaling with 3D reference points provided by structure-from-motion. Tab. III and Tab. IV provide zero-shot performance on indoor and outdoor datasets, respectively. For each dataset, our results show rescaling using LiDARs with different numbers of beams. Additionally, we present results of rescaling with structure-from-motion for the DDAD dataset [guizilini20203d]. We do not provide results of rescaling with structure-from-motion on NYUv2 [nyu], SUN-RGBD [song2015sun], IBIMS-1 [ibims] and DIODE Indoor and Outdoor [diode_dataset] because they do not consist of temporal image sequences with poses. The results of the monocular depth estimation methods we compare against were directly taken from their respective papers, which explain missing numbers when they have not been reported by their authors.
@@
Our experiments highlight an overall benefit of using our rescaling approach with $32$-laser or $16$-laser LiDAR rather than using zero-shot monocular metric depth estimation. Thus, we observe on average 4646% improvement on $\delta_{1}$, 0.50.5% on AbsRel and 4747% on RMS relative to the other zero-shot methods. In contrast, the methods that have been fine-tuned on the same domain as the test set (see lines with (FT) in Tab. I and Tab. II) compare favorably to ours. These approaches benefit from an additional in-domain training which is often costly due to dataset creation and training computation. We notice that our method is sensitive to the domain with on average 2828% and 1313% enhancements on indoor (including NYUv2) and outdoor (including KITTI) datasets, respectively, relative to other zero-shot methods whatever the metric. Regarding performance with 2D LiDAR, the same sensitivity to the domain is apparent. However, the results are a bit lower than zero-shot and fine-tuned metric depth estimation methods and lower than our other approaches, especially on outdoor datasets while remaining competitive on indoor datasets. Interestingly, we observe that increasing the number of beams from 16 to 32 does not necessarily improve the performance.